# biblio.bib

### v2

@@ Abstract
Many organisms and cell types, from bacteria to cancer cells, exhibit a remarkable ability to adapt to fluctuating environments. Additionally, cells can leverage memory of past environments to better survive previously-encountered stressors. From a control perspective, this adaptability poses significant challenges in driving cell populations toward extinction, and is thus an open question with great clinical significance. In this work, we focus on drug dosing in cell populations exhibiting phenotypic plasticity. For specific dynamical models switching between resistant and susceptible states, exact solutions are known. However, when the underlying system parameters are unknown, and for complex memory-based systems, obtaining the optimal solution is currently intractable. To address this challenge, we apply reinforcement learning (RL) to identify informed dosing strategies to control cell populations evolving under novel non-Markovian dynamics. We find that model-free deep RL is able to recover exact solutions and control cell populations even in the presence of long-range temporal dynamics.
@@ 1 Introduction
In order to survive, organisms must adapt to unpredictable environmental stressors occurring over diverse timescales. As a result, biological systems display remarkable adaptive capabilities, making it exceptionally challenging to control them through environmental modulation—an open and significant question in the physics of living systems. Two examples of adaptive systems with great clinical importance are cancer cell resistance to chemotherapy [Vasan2019], and bacterial resistance to antibiotics [Blair2015]. In both settings, constant application of a drug does not typically result in population extinction, as drug application also drives a certain fraction of the population to alter their phenotypic state to become drug-resistant. This new resistant state can persist even after drug removal and across cell lineages, thus encoding a memory of the stressful environment which allows populations to more quickly adapt if the drug is reapplied [Harmange2023, Mathis2017, Banerjee2021a].
Previous work [padmanabhan2017reinforcement, engelhardt2020dynamic, Gallagher2024, fischer2024minimising] has studied temporal drug dosing protocols to slow or prevent adaptation in various cancer or bacterial models using different methods. However, these models fail to account for the variety of adaptive timescales present in real biological systems. The presence of such memory effects greatly complicates the control problem, making the study of more realistic models imperative. Thus, to study the control of such memory-based adaptive systems, in this work we introduce a novel population model exhibiting phenotypic plasticity with non-Markovian dynamics. We couple insights from control theory with deep reinforcement learning to discover interpretable treatment protocols which successfully prevent proliferation.
@@ 2 Proposed Model and Approach
Refer to caption Refer to caption Refer to caption
Figure 1: Left: Interaction loop between RL agent and environment (Sec. 2.2). Middle: depiction of the phenotypic switching model (Sec. 2.1). Right: Effect of learned policy on resistant fraction. for different memory strengths.
@@ 2.1 Non-Markovian Phenotypic Switching Model
To model the treatment response of an adaptive cell population, we use a general phenotypic switching model which captures the time evolution of a susceptible subpopulation, with size $S(t)$, and a resistant subpopulation, with size $R(t)$. Phenotypic switching models have been successful in describing a wide variety of biological scenarios, including development of persister cells and resistant cells in bacteria, and development of drug resistance in cancer cells [Balaban2004, fischer2024minimising, Witzany2023, kumar2019stochastic, Kratz2024]. In our model, susceptible cells with a net growth rate $\kappa_{S}(u)$ ($\kappa_{S}(u)<0$ corresponds to net cell death) switch to a resistant state at a rate $\alpha(u)$, where $u$ is the drug concentration. Similarly, when the drug is removed, resistant cells with a net growth rate $\kappa_{R}(u)$ switch back to the susceptible state at a concentration-dependent rate $\delta(u)$ (Fig. 1, middle). All growth and switching rates are a function of drug concentration normalized by the maximum allowable dose, thus $u\in[0,1]$. The time evolution of the size of each subpopulation, $\textbf{x}(t)=[S(t),R(t)]^{T}$, is then given by the dynamical system:
$\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t),u(t))=\textbf{A}(u(t))\textbf{x}(t)$ (1)
@@
We choose $\kappa_{S}(u)$ and $\kappa_{R}(u)$ to decrease and increase with drug concentration, respectively. This parametrization corresponds to “drug addiction” behavior in the resistant subpopulation, a robust phenomenon which has been observed not only in cell culture [SUDA2012292, Sun2014, Moriceau2015], but in animal models [DasThakur2013] and in vivo [Seifert2016138, Dooley2016]. Similarly, $\delta(u)$ decreases with $u$ while $\alpha(u)$ increases with $u$, as drug application drives the population to become more resistant, while reduction in drug concentration causes the system to recover susceptibility (see Appendix A for complete model details).
Recently, cell populations of many types, including human cancer cell lines, yeast, and bacteria, have been shown to maintain a memory of past environments which facilitates adaptation to previously-seen stressors over many timescales [Shaffer2020, Harmange2023, Larkin2024, Wolf2008, Mathis2017]. To better capture this memory dependence on treatment response, we introduce a memory kernel into the previously-described dynamics (Eq. (1)), making them non-local in time. Specifically, we choose a fractional differential equation (FDE) formulation as a phenomenological way to introduce multiple timescales of adaptation, one which has been used successfully to model memory effects in other biological [Lundstrom2008], ecological [Khalighi2022], and physical contexts [Bonfanti2020]. With this addition the dynamics now become:
@@
where $\Gamma(\cdot)$ denotes the Gamma function, $\textbf{f}(\cdot)$ is given by Eq. (1), and here we introduce the parameter $\mu\in(0,1]$ which controls memory strength. A value of $\mu=1$ corresponds to the memoryless case (first order derivative), whereas smaller values of $\mu$ correspond to an increased influence of past states on the current dynamics (lower order fractional derivative).
We seek to obtain a temporal drug concentration protocol $u(t)$ which minimizes the growth of a population and thus choose the final cost as $C:=\log N(T)/N(0)$ over the time interval $T$, where $N(t)=S(t)+R(t)$ represents the total population. Thus, we aim to solve the following control problem:
@@
where $\textbf{x}(T;u(\cdot))$ denotes the state of x at terminal time $T$ subject to control $u$ from $0\leq t\leq T$. Interestingly, in our minimal model (Eq. (3)), as long as $\kappa_{S}(u)$, $\kappa_{R}(u)$, $\alpha(u)$, and $\delta(u)$ are monotonic functions, we can show that the optimal control solution follows “bang-bang” control, regardless of model parameters and memory strength (see Appendix A for details). Thus, the continuous state-transition matrix of Eq. (2) can be simplified into a discrete one with binary control inputs, without altering the optimal solution (Appendix A). Despite this simplification, model control remains difficult, as the number of times and duration of drug application must be optimized. Constant application of the drug at the maximum dose ($u=1$) results in cell adaptation and proliferation (Fig. 2): a highly suboptimal solution to Problem (4) and a catastrophic result in the clinical context. Previous work [fischer2024minimising] has shown that in the memoryless case ($\mu=1$), the optimal solution for this type of model requires an initial drug application phase, followed by pulsing between treatment and pause phases at a regular interval dependent on the model parameters. However, as seen in Fig. 2, we find that the addition of memory ($\mu<1$) renders the control strategy of the memoryless case ineffective, as cells which have previously encountered treatment switch to the resistant state faster upon subsequent applications. Furthermore, in clinical or experimental setting, obtaining the values which parameterize Eq. (3) is usually not feasible, thus obtaining the appropriate switching frequency through direct computation via optimal control (OC) theory becomes impossible. As a result, we seek to learn the optimal policy directly through experience using reinforcement learning.
2.2 Reinforcement Learning
Reinforcement learning (RL) allows an agent to learn a drug protocol through experience (Fig. 1, left), despite the non-Markovian dynamics and without access to the underlying environment-specific model parameters. To formulate the RL problem, we define the relevant characteristics as follows: State: The state ($s_{t}$) of the agent’s environment is a list of the last $K=5$ estimates of the instantaneous growth rate, $c_{t}={\Delta^{-1}}\log N_{t}/N_{t-\Delta}$, where $\Delta$ is the simulation time between actions. Thus, a history of past observations is encoded in the state vector, a crucial design choice if the agent is to learn control without a recurrent hidden state. Action: As motivated in 2.1, we choose a binary action space $u\in\{0,1\}$ representing whether the drug is applied or not, as this is suitable to recover optimal control. Reward: As we seek to solve Problem (4), the reward is simply the negative growth rate, $r_{t}=-c_{t}$. Notice that with this choice of reward function, the sum of rewards across a trajectory simplifies as $R_{0:T}=\Delta^{-1}\log N_{0}/N_{T}$, ensuring the agent’s objective is aligned with a reduction in total cell population. Dynamics: The dynamics of the total cell population $N_{t}$ are governed by the dynamical system described in 2.1, initialized to be fully susceptible ($\textbf{x}_{0}=[1000,0]$). After an action is executed, the simulation (a numerical solution  [^1] The numerical solution of fractional differential equations requires some care; cf. [garrappa2018numerical] for details. of Eq. (3)) is computed with the action (dose) fixed for $\Delta=0.01$ hours to compute the next state ($s_{t+\Delta}$).
We use our own implementation of Double DQN [van2016deep] based on open-source code for DQN [stable-baselines3] to train the agent in this environment. After an action is taken, it is stored in an experience replay buffer for later use. We parameterize the value function $Q$ with a neural network (with parameters $\theta$), and train via SGD by sampling mini-batches uniformly at random from the buffer. The loss function is defined as the Bellman residual loss – the squared difference between left and right hand sides of the following equation:
As common in value-based algorithms, we use an additional target network and use exponentially annealed $\varepsilon$-greedy exploration. Further details on reinforcement learning can be found e.g. in [sutton&barto, rainbow]. We tune over several hyperparameters (whose values we list in the Appendix) to optimize performance. To encourage the agent to reduce the cell population while decreasing runtime, we terminate the episode if the number of cells ever exceeds its initial amount, forcing the agent to initially apply a continuous dose. All code to reproduce our experimental results can be found at https://github.com/JacobHA/RL4Dosing.
3 Results
Figure 2: Performance comparison of constant drug application, solution for the memoryless case, resistant fraction-based pulsing technique, and policy learned by RL. For the fraction-based policy, an optimal lower and upper bound for resistant fractions are found through sweeping (Appendix A.1). The RL policy is capable of controlling the cell population better than any other scheme.
We first test DQN in the memoryless case ($\mu=1$), for which an optimal controller is known [fischer2024minimising]. In this case, the optimal policy can be derived based only on two consecutive resistant fractions. However, given only the growth rates, the RL agent is able to reliably recover the optimal policy with only two frames and without any access to underlying model parameters. It does so by first applying the drug and then pulsing regularly, in agreement with the OC solution, as seen in Fig. 2 (rightmost plot). With confirmation that RL can find the optimal dosing strategy in the memoryless case, we turn to the more difficult memory-based dynamics ($\mu<1$). We do not have a solution to Problem 4 in this regime, so we compare to two baselines: The memoryless protocol and a modified version in which switching times occur when the resistant fraction reaches a threshold value. We note that these baselines can have arbitrarily small switching times, which is practically infeasible. Remarkably, we find that using a small but bounded $\Delta$, our learned policy outperforms both of these baselines (Fig. 2). In addition, our experiments show that decreasing $\Delta$ beyond a certain threshold does not considerably increase performance (Fig. 4). Interestingly, the learned policy reveals that the agent initially maintains constant drug application before transitioning to a memory-dependent pulsing protocol. The agent increases the frequency of pulsing throughout the trajectory until saturation at the maximum rate ($\Delta^{-1}$) (Fig. 3). For cases with memory, this increase in dosing frequency can be understood as the result of faster cellular adaptation back to the resistant state after multiple drug encounters. Thus, to maintain a susceptible population, the policy’s pulse frequency must continuously be increased to compensate for the memory-based adaptability. We also find that the agent maintains a lower average resistant fraction at higher memory values (Fig. 1, right).
4 Discussion
In this work, we study the control of a highly non-Markovian model of adaptive cellular growth dynamics using deep reinforcement learning. Although we focus here on a specific parameterization most relevant to cancer, we expect this methodology to be applied successfully to other scenarios, including resistance development in bacteria. We find that deep RL is capable of recovering the known optimal policy for the memoryless case and can successfully find a policy for memory-based systems which prevents proliferation, all without access to the underlying model parameters. We utilize frame-stacking [nature-dqn] to encapsulate the history of the agent’s trajectory, but in future work we will test the use of recurrent policies to capture more nuanced long-term effects and perhaps further improve performance. We also plan to extend our deterministic framework to the stochastic setting, as population heterogeneity is known to further complicate the control process. This work demonstrates the benefits of combining OC and RL, showing how their frameworks can be effectively integrated. This connection can be further developed by using the memoryless OC solution to enhance RL training through reward shaping [ng1999policy, adamczyk2023utilizing] or pre-training [uchendu2023jump] techniques.
Experimentally measuring the growth rate of a pathogenic population is easier than determining its resistant fraction. Surprisingly, we find that RL can learn successful policies directly from the growth rate making it a promising method for use in clinical settings.
5 Acknowledgements
JA acknowledges funding support from the Topical Group on Data Science (GDS) of the American Physical Society; the NSF through Award No. PHY-2425180; the use of the supercomputing facilities managed by the Research Computing Department at UMass Boston; and fruitful discussions with Rahul V. Kulkarni and Stas Tiomkin. JCK acknowledges support from the NeurIPS ML4PS organizers, the Computational Biology Department at CMU, and from Shiladitya Banerjee in the Department of Physics at CMU. This project developed out of the hackathon at the 2024 Summer School hosted by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI).
@@ 5 Acknowledgements
Appendix A Optimal control of the phenotypic switching model yields a bang-bang control solution.
We seek to find a temporal drug protocol $u(t)$ which solves the control problem
$\min_{u}C(\textbf{x}(0),\textbf{x}(T;u(\cdot)))\;\;\text{subject to}\;\;\dot{\textbf{x}}=\textit{{F}}(\textbf{x}(t),u(t)),\;\textbf{x}(0)=\textbf{x}_{0},\;u(t)\in[0,1].$ (6)
We choose the terminal cost to be the logarithm of the relative growth of the population throughout the course of treatment, $C(\textbf{x}(0),\textbf{x}(T;u(\cdot))):=\log N_{T}/N_{0}$, where $\textbf{x}(T;u(\cdot))$ denotes the state of x at terminal time $T$ subject to control $u$ from $0\leq t\leq T$, and where $N(t)=S(t)+R(t)$. Using Eq. (1) and the fact that $N(t)=S(t)+R(t)$, the model dynamics can be rewritten as a single fractional differential equation, namely:
where $D_{0}^{\mu}$ denotes the Caputo fractional derivative of order $\mu$ starting at $t=0$ [garrappa2018numerical], and here we drop the explicit time dependence for notational clarity. This can equivalently be written as the continuous delay differential equation (we use this form in the main text):
[…]
$\kappa_{S}(u)=\kappa_{S}^{\mathrm{max}}-(\kappa_{S}^{\mathrm{max}}-\kappa_{S}^{\mathrm{min}})g(u)\;,\;\;\;\kappa_{R}(u)=\kappa_{R}^{\mathrm{min}}+(\kappa_{R}^{\mathrm{max}}-\kappa_{R}^{\mathrm{min}})g(u)\;,$ (9)
where $g(u)\in[0,1]$ is a monotonic dose-response function which relates drug dose to net growth rate, $\kappa_{S}^{\mathrm{max}}>0$ denotes the maximum growth rate of the susceptible subpopulation in the absence of drug application, and where $\kappa_{S}^{\mathrm{min}}<0$ is the maximum death rate of the susceptible subpopulation caused by application of the maximum drug dose ($u=1$). Importantly, $\kappa_{R}^{\mathrm{max}}>0$ denotes the maximum growth rate of the resistant subpopulation, which occurs in the presence of the maximum drug dose, and $\kappa_{R}^{\mathrm{min}}<0$ corresponds to the maximum death rate of the resistant subpopulation, which occurs when the drug is removed. This parametrization corresponds to “drug addiction” behavior, a phenomenon observed in several types of cancers [SUDA2012292, Sun2014, Moriceau2015, Seifert2016138, DasThakur2013, Dooley2016]. Similarly, the switching rates can then be defined as:
$\alpha(u)=\alpha_{\mathrm{max}}g(u)\;\;\text{and}\;\;\delta(u)=\delta_{\mathrm{max}}(1-g(u))\;,$ (10)
where $\alpha_{\mathrm{max}},\delta_{\mathrm{max}}>0$ denote the phenotypic switching rates of cells switching from the susceptible state to the resistant state, and vice versa.
@@
where the trajectory of the Lagrangian multiplier $\lambda(t)$ is the solution to the costate equation [Gomoyunov2023]:
@@
Applying Pontryagin’s minimum principle, we obtain the resulting inequality
@@
which along with Eqs. (7) and (11) can be used to obtain the optimal control policy:
$u^{*}(t)=\arg\min_{u}g(u(t))B(\phi^{*}(t),\lambda^{*}(t))\;,$ (14)
where $B(\phi,\lambda)=\lambda\alpha_{\mathrm{min}}-\lambda(\kappa_{S}^{\mathrm{max}}-\kappa_{S}^{\mathrm{min}}+\kappa_{R}^{\mathrm{max}}-\kappa_{R}^{\mathrm{min}})\phi^{2}+\lambda(\kappa_{R}^{\mathrm{max}}-\kappa_{R}^{\mathrm{min}}+\kappa_{S}^{\mathrm{max}}-\kappa_{S}^{\mathrm{min}}+\delta_{\mathrm{min}}-\alpha_{\mathrm{min}})\phi$. As long as $g(u)$ is a monotonically increasing function of $u$, then the resulting control is said to be “bang-bang”, where $u(t)$ only takes extreme values. The switching times between maximum and minimum values of $u$ is determined by $B(\phi(t),\lambda(t))$, the switching function, yielding the optimal control solution:
$u^{*}(t)$ $=0\;\;\text{if}\;\;B(\phi^{*}(t),\lambda^{*}(t))>0,$
$u^{*}(t)$ $=1\;\;\text{if}\;\;B(\phi^{*}(t),\lambda^{*}(t))<0,\;\;\text{and}$
@@
In principle, the optimal control trajectory can be obtained through numerical integration of the model dynamics (Eq. (7)) along with the corresponding costate equation (Eq. (12)). In the context of bang-bang control on non-Markovian systems however, using this approach can be difficult, as it requires careful choice of integration technique and update rule. Thus, we turn to reinforcement learning.
As shown above, the optimal control solution to Problem (6) follows bang-bang control, regardless of model parameters. This allows the continuous model of Eq. (1) to be simplified to a discrete model with binary controls without altering the optimal solution. This yields:
@@
A.1 Obtaining the optimal solution in the memoryless case

### v3

# Reinforcement Learning for Control of Non-Markovian Cellular Population Dynamics

@@ Abstract
Many organisms and cell types, from bacteria to cancer cells, exhibit a remarkable ability to adapt to fluctuating environments. Additionally, cells can leverage a memory of past environments to better survive previously-encountered stressors. From a control perspective, this adaptability poses significant challenges in driving cell populations toward extinction, and thus poses an open question with great clinical significance. In this work, we focus on drug dosing in cell populations exhibiting phenotypic plasticity. For specific dynamical models switching between resistant and susceptible states, exact solutions are known. However, when the underlying system parameters are unknown, and for complex memory-based systems, obtaining the optimal solution is currently intractable. To address this challenge, we apply reinforcement learning (RL) to identify informed dosing strategies to control cell populations evolving under novel non-Markovian dynamics. We find that model-free deep RL is able to recover exact solutions and control cell populations even in the presence of long-range temporal dynamics. To further test our approach in more realistic settings, we demonstrate robust RL-based control strategies in environments with measurement noise and dynamic memory strength.
@@ 1 Introduction
In order to survive, organisms must adapt to unpredictable environmental stressors occurring over diverse timescales. As a result, biological systems display remarkable adaptive capabilities, making it exceptionally challenging to control them through environmental modulation, marking an open and significant question in the physics of living systems. Two examples of adaptive systems with great clinical importance are cancer cell resistance to chemotherapy (Vasan et al., 2019), and bacterial resistance to antibiotics (Blair et al., 2015). In both settings, the control problem of drug application is complicated by unknown model parameters and non-Markovian dynamics. Notably, the constant application of a drug does not typically result in population extinction, as drug application also drives a certain fraction of the population to alter their phenotypic state to become drug-resistant. This new resistant state can then persist even after drug removal and across cell lineages, thus encoding a memory of the stressful environment which allows populations to more quickly adapt if the drug is reapplied Harmange et al. (2023); Mathis & Ackermann (2017); Banerjee et al. (2021). As a result, it is unclear how to devise a control strategy that can effectively mitigate long-term population growth in these scenarios. We address this open problem in the present work.
Previous literature (Padmanabhan et al., 2017; Engelhardt, 2020; Gallagher et al., 2024; Fischer & Bluethgen, 2024) has studied temporal drug dosing protocols to slow or prevent adaptation in various cancer or bacterial models using various methods. However, these models fail to account for the variety of adaptive timescales present in real biological systems. The presence of such memory effects greatly complicates the control problem, making the study of more realistic models imperative. Thus, to study the control of such memory-based adaptive systems, in this work we introduce a novel population model exhibiting phenotypic plasticity with non-Markovian dynamics. Prior work has shown that learning-based methods are able to form the basis of patient-specific treatment protocols (Miotto et al., 2016; Ahn et al., 2021). We show that insights from control theory can be coupled with deep reinforcement learning to discover treatment protocols which successfully prevent proliferation. We highlight our main contributions below:
Main Contributions
  1. 1.
    We develop a novel memory-based model for phenotypic switching relevant for bacterial and cancerous population dynamics.
  2. 2.
    We pose the problem of optimal drug dosing in this model and provide insights on the solution from control theory.
  3. 3.
    We find a robust optimal policy with deep RL that is independent of memory strength, requiring only clinically-accessible observations.
In the following sections, we first introduce the new population model, derive bang-bang optimal control, and experimentally show that bang-bang control is required to find a high-performing policy. Because of the non-Markovian dynamics, we introduce a short history to the agent’s state which we implement via framestacking. Using ideas from optimal control theory and deep reinforcement learning, we thus simplify the original control problem in a well-principled manner. We then apply standard and more advanced RL techniques to solve different versions of the posed problem.
@@ 2 Proposed Model and Approach
In this section, we develop our novel non-Markovian model for cellular dynamics. These dynamics are based on a simple but general switching model (Figure 1) with switching rates that are dependent on drug concentration. We then introduce a physically motivated non-linear memory kernel, before discussing general approaches to the optimal control problem.
@@ 2.1 Non-Markovian Phenotypic Switching Model
Figure 1: Depiction of the deterministic phenotypic switching model. The susceptible subpopulation $S$ transitions to the resistant state $R$ at a concentration-dependent rate, $\delta$. Similarly, the resistant state switches back to a susceptible state at a rate $\alpha$. The subpopulations each have a concentration-dependent growth or death rate, $\kappa$.
To model the treatment response of an adaptive cell population, we use a general phenotypic switching model which captures the time evolution of a susceptible subpopulation, with size $S(t)$, and a resistant subpopulation, with size $R(t)$. (The notation for susceptible and resistant states should not be confused with state and reward in RL, for which we use lowercase letters.) Phenotypic switching models have been successful in describing a wide variety of biological scenarios, including development of persister cells and resistant cells in bacteria, and development of drug resistance in cancer cells (Balaban et al., 2004; Fischer & Bluethgen, 2024; Witzany et al., 2023; Kumar et al., 2019; Kratz & Banerjee, 2024). In our model, susceptible cells with a net growth rate $\kappa_{S}(u)$ ($\kappa_{S}(u)<0$ corresponds to net cell death) switch to a resistant state at a rate $\alpha(u)$, where $u$ is the drug concentration. Similarly, when the drug is removed, resistant cells with a net growth rate $\kappa_{R}(u)$ switch back to the susceptible state at a concentration-dependent rate $\delta(u)$ (Fig. 1). All growth and switching rates are a function of drug concentration normalized by the maximum allowable dose, thus $u\in[0,1]$. The time evolution of the size of each subpopulation, $\textbf{x}(t)=[S(t),R(t)]^{T}$, is then given by the dynamical system:
$\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t),u(t))=\textbf{A}(u(t))\textbf{x}(t)\;,$ (1)
@@
$\kappa_{S}(u)=\kappa_{S}^{\mathrm{max}}-(\kappa_{S}^{\mathrm{max}}-\kappa_{S}^{\mathrm{min}})g(u)\;,\;\;\;\kappa_{R}(u)=\kappa_{R}^{\mathrm{min}}+(\kappa_{R}^{\mathrm{max}}-\kappa_{R}^{\mathrm{min}})h(u)\;,$
where $g(u),h(u)\in[0,1]$ is a monotonic dose-response function which relates drug dose to net growth rate, $\kappa_{S}^{\mathrm{max}}>0$ denotes the maximum growth rate of the susceptible subpopulation in the absence of drug application, and where $\kappa_{S}^{\mathrm{min}}<0$ is the maximum death rate of the susceptible subpopulation caused by application of the maximum drug dose ($u=1$). Importantly, $\kappa_{R}^{\mathrm{max}}>0$ denotes the maximum growth rate of the resistant subpopulation, which occurs in the presence of the maximum drug dose, and $\kappa_{R}^{\mathrm{min}}<0$ corresponds to the maximum death rate of the resistant subpopulation, which occurs when the drug is removed. This parametrization corresponds to “drug addiction” behavior in the resistant subpopulation, a robust phenomenon which has been observed not only in cell culture (Suda et al., 2012; Sun et al., 2014; Moriceau et al., 2015), but in animal models (Das Thakur et al., 2013) and in vivo (Seifert et al., 2016; Dooley et al., 2016). Similarly, $\delta(u)$ decreases with $u$ while $\alpha(u)$ increases with $u$, as drug application drives the population to become more resistant, while reduction in drug concentration causes the system to recover susceptibility. These switching rates can then be defined as:
$\alpha(u)=\alpha_{\mathrm{max}}j(u)\;\;\text{and}\;\;\delta(u)=\delta_{\mathrm{max}}(1-k(u))\;,$
where $j(u),k(u)\in[0,1]$ and where $\alpha_{\mathrm{max}},\delta_{\mathrm{max}}>0$ denote the phenotypic switching rates of cells switching from the susceptible state to the resistant state, and vice versa.
Recently, cell populations of many types, including human cancer cell lines, yeast, and bacteria, have been shown to maintain a memory of past environments which facilitates adaptation to previously-seen stressors over many timescales (Shaffer et al., 2020; Harmange et al., 2023; Larkin et al., 2024; Wolf et al., 2008; Mathis & Ackermann, 2017). To better capture this memory dependence on treatment response, we introduce a memory kernel into the previously-described dynamics (equation 1), making them non-local in time. Specifically, we choose a fractional differential equation (FDE) formulation as a phenomenological way to introduce multiple timescales of adaptation, one which has been used successfully to model memory effects in other biological (Lundstrom et al., 2008), ecological (Khalighi et al., 2022), and physical contexts (Bonfanti et al., 2020). With this addition the dynamics now become:
@@
where $\Gamma(\cdot)$ denotes the Gamma function, $\textbf{f}(\cdot)$ is given by equation 1, and here we introduce the parameter $\mu\in(0,1]$ which controls memory strength. A value of $\mu=1$ corresponds to the memoryless case (first order derivative), whereas smaller values of $\mu$ correspond to an increased influence of past states on the current dynamics (lower order fractional derivative).
We seek to obtain a temporal drug concentration protocol $u(t)$ which minimizes the growth of a population and thus choose the final cost as $C:=\log N(T)/N(0)$ over the time interval $T$, where $N(t)=S(t)+R(t)$ represents the total population.
Thus, we aim to solve the following control problem:
@@
where $\textbf{x}(T;u(\cdot))$ denotes the state of x at terminal time $T$ subject to control $u$ from $0\leq t\leq T$. Interestingly, in our minimal model (equation 3), as long as $\kappa_{S}(u)$, $\kappa_{R}(u)$, $\alpha(u)$, and $\delta(u)$ are monotonic functions, we can show that the optimal control solution follows “bang-bang” control, regardless of the precise model parameters and memory strength.
2.2 Derivation of bang-bang control
Using equation 1 and the fact that $N(t)=S(t)+R(t)$, the model dynamics can be rewritten as a single fractional differential equation in terms of the fraction of resistant cells, $\phi\coloneqq R/N$, namely:
where $D_{0}^{\mu}$ denotes the Caputo fractional derivative of order $\mu$ starting at $t=0$ (Garrappa, 2018), and here we drop the explicit time dependence for notational clarity. This can equivalently be written as the continuous delay differential equation (we use this form in equation 3):
@@ 2.2 Derivation of bang-bang control
where the trajectory of the Lagrangian multiplier $\lambda(t)$ is the solution to the costate equation (Gomoyunov, 2023):
@@
Applying Pontryagin’s minimum principle, we obtain the following inequality
@@
which along with equation 5 and equation 7 can be used to obtain the optimal control policy:
$u^{*}=\arg\min_{u}\lambda^{*}[\Delta_{S}(\phi^{*}-\phi^{*2})g(u)+\Delta_{R}(\phi^{*}-\phi^{*2})h(u)+\delta_{\mathrm{max}}\phi^{*}k(u)+\alpha_{\mathrm{max}}(1-\phi^{*})j(u)]\;,$
where the following shorthand notation is introduced: $\Delta_{S}=\kappa_{S}^{\mathrm{max}}-\kappa_{S}^{\mathrm{min}}$ and $\Delta_{R}=\kappa_{R}^{\mathrm{max}}-\kappa_{R}^{\mathrm{min}}$.
Critically, the fact that $\kappa_{S}^{\mathrm{max}},\kappa_{R}^{\mathrm{max}},\alpha_{\mathrm{max}},\delta_{\mathrm{max}}>0$, $\kappa_{S}^{\mathrm{min}},\kappa_{R}^{\mathrm{min}}<0$, and $g,\ h,\ j,\ k,\ \phi\in[0,1]$ ensures that the bracketed sum which multiplies $\lambda^{*}$ is positive for $u>0$. As a result, as long as $g(u),\ h(u),\ j(u)$, and $k(u)$ are monotonically increasing functions of $u$, then the resulting optimal control is said to be “bang-bang”, where $u^{*}(t)$ only takes on its extreme values. Furthermore, the switching times between the maximum and minimum values of $u$ are determined by $\lambda^{*}(t)$, yielding the optimal control solution:
$u^{*}(t)$ $=0\;\;\text{if}\;\;\lambda^{*}(t)>0,$
$u^{*}(t)$ $=1\;\;\text{if}\;\;\lambda^{*}(t))<0,\;\;\text{and}$
@@
In principle, the optimal control trajectory can be obtained through numerical integration of the model dynamics (equation 5) along with the corresponding costate equation (equation 8). In the context of bang-bang control of non-Markovian systems however, using this approach can be difficult, as it requires careful choice of integration technique and update rule. Thus, we turn to reinforcement learning.
As shown above, the optimal control solution to Problem 4 follows bang-bang control, regardless of model parameters. This allows the continuous model of equation 1 to be simplified to a discrete model with binary controls without altering the optimal solution. This yields:
@@
2.3 Control Strategies
Despite this simplification in action space (from continuous to binary controls), the optimal control problem remains difficult, as the number of times and duration of drug application must be optimized. Constant application of the drug at the maximum dose ($u=1$) results in cell adaptation and proliferation (Fig. 3): a highly suboptimal solution to Problem 4 and a catastrophic result in the clinical context.
Recently,(Fischer & Bluethgen, 2024) has shown that in the memoryless case ($\mu=1$), the optimal solution for this type of model requires an initial drug application phase, followed by pulsing between treatment and pause phases at a regular interval dependent on the model parameters. However, as seen in Fig. 3, we find that the addition of memory ($\mu<1$) renders the control strategy of the memoryless case ineffective, as cells which have previously encountered treatment switch to the resistant state faster upon subsequent applications. Furthermore, in the clinical or experimental setting, obtaining the values which parameterize equation 3 is usually not feasible, and one can only rely on direct and more macroscopic measurements from the cell populations. Thus, obtaining the appropriate switching frequencies through direct computation via optimal control (OC) theory becomes impossible. One may propose to learn these values directly before feeding them into an OC solution, but this may result in a brittle pipeline, especially as the model discussed may not fully describe the underlying biological dynamics. As a result, we seek to learn the optimal policy end-to-end, directly through experience using deep reinforcement learning (Sec. 3).
2.4 Obtaining the optimal solution in the memoryless case